OpenStudy (anonymous):

Please help me step by step!!! ASAP!!!!! 2. evaluate the determinant of the matrix {-4 5 6] [0 4 4] [-2 -5 4]

OpenStudy (rock_mit182):

\[-4\left[\begin{matrix}4 & 4 \\ -5 & 4\end{matrix}\right]-5\left[\begin{matrix}0 & 4 \\ -2 & 4\end{matrix}\right]+6\left[\begin{matrix}0 & 4 \\ -2 & -5\end{matrix}\right]\]

OpenStudy (anonymous):

okay so wait the first one you multiply all of it by -4 right ??

OpenStudy (rock_mit182):

yes, and where did come from that -4

OpenStudy (amistre64):

i do the shortcut ofr a 3x3 ...

OpenStudy (rock_mit182):